In a class there are 22 girls and 38 boys. The mean score of girls in an examination is 78 and the mean score of boys is 71. What is the mean score of the entire class ?
Answer by srinivas.g(540)

sum of scores of all boys = no of boys * mean score of boys
= 38*71
=2698
sum of scores of all girls = no of girls * mean score of girls
= 22*78
=1716
total =2698+1716 =4414
mean = total / total no of boys girls
=4414/(22+38)
=4414/60
=73.567
mean score of entire class = 73.567